OpenStudy (anonymous):

use the elimination method to solve the system of equations. x+4y=37 -x+3y=19

HanAkoSolo (jamierox4ev3r):

hello there Avon! x and -x cancel each other out, so you are left with 4y=37 and 3y=19

HanAkoSolo (jamierox4ev3r):

Do you think you can solve these two equations now? @Avon ?

OpenStudy (anonymous):

@Jamierox4ev3r 7y=56

HanAkoSolo (jamierox4ev3r):

right, then divide 56 by 7. What do you get? @avon?

OpenStudy (anonymous):

8 @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

right, so that means that y=8.

HanAkoSolo (jamierox4ev3r):

THen, plug in the y value (8) into both given equations and solve for x. Try to do that on your own but if you don't understand, I will help you okay?

OpenStudy (anonymous):

idk to use the top or bottom equation 2 solve for x @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

you use both equations! and if they both equal the same thing, then you are correct ^_^

HanAkoSolo (jamierox4ev3r):

does that make sense @Avon ?

OpenStudy (anonymous):

no

HanAkoSolo (jamierox4ev3r):

let me use one as an example x+4(8)=37 x+32 =37 -32 -32 --------- x=5

HanAkoSolo (jamierox4ev3r):

for the first equation, x equals 5. Now , do the second equation and if it also equals 5, that means you are correct!

OpenStudy (anonymous):

o ok i c now

HanAkoSolo (jamierox4ev3r):

:) cool

OpenStudy (anonymous):

so (8,5) or the solution set

HanAkoSolo (jamierox4ev3r):

actually it is (5,8) because the x variable always is listed before the y variable. However, have you tried the second equation yet?

HanAkoSolo (jamierox4ev3r):

but yeah that is correct! nice job @Avon